OpenStudy (anonymous):

What is the volume of a cube with surface area 54 times (x ^ 2)?

OpenStudy (anonymous):

Surface area of a cube with side length \(s\) is \(6s^2\). \[6s^2=54x^2~~\iff~~s^2=9x^2~~\iff~~s=3x\] Volume of a cube with the same side length is \(s^3\), so \[V=(3x)^3=27x^3\]